What is. Moment of inertia of a solid hemisphere and a solid cone?

Grade:11

1 Answers

Arun
6 years ago
Dear Tarun
 
Momentum of inertia of hemisphere about the axis shown in fig. is I=(1/5)MR^2. But M = 2m, where, m is mass of hemisphere, then,
I = (2/5)mR^2.
Moment of inertia of solid cone= 3/10*m*r^2

Regards
Arun (askIITians forum expert)